OpenStudy (anonymous):

What is the value of 8C5? A. 40 B. 56 C. 960 D. 6720

OpenStudy (anonymous):

i will fan and medal

Nnesha (nnesha):

\[\huge\rm \frac{ n! }{ r!(n-r)! }\] formula for combination n_C_r n=8 r=5

OpenStudy (anonymous):

huh

Nnesha (nnesha):

substitute n for 8 and r for 5

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

just tell me the answer

Nnesha (nnesha):

use calculator replace variables with their values that's it!

OpenStudy (anonymous):

im not good at it help me

Nnesha (nnesha):

[Math Processing Error]\[\huge\rm \frac{ 8! }{ 5!(8-5)! }\] formula for combination n_C_r n=8 r=5 can you solve now ?

OpenStudy (anonymous):

c

Nnesha (nnesha):

no..

OpenStudy (anonymous):

a

Nnesha (nnesha):

nice try now stop guessing :)

Nnesha (nnesha):

CAN you use a calculator ??

OpenStudy (anonymous):

i wasnt guessing those are what i got

Nnesha (nnesha):

how did you get A and C show your work then.

OpenStudy (anonymous):

is it b

Nnesha (nnesha):

or just do it b hands solve the parentheses 8-5 = 3! \[\huge\rm \frac{ 8! }{ 5!(8-5)! }\] formula for combination n_C_r n=8 r=5 \[\frac{ 8! }{ 5!(3!) }\] 8!= 8 times 7 times 6 times 5 times 4 times 3 times 2 times 1 subtract one untill u get one

Nnesha (nnesha):

or 8!= 8 times 7 times 6 times 5! <--now you can divide 5! \[\huge\rm \frac{ 8 \times 7 \times 6 \times\cancel{ 5!} }{\cancel{ 5!} 3! }\] now 3!= ??

OpenStudy (anonymous):

d

Nnesha (nnesha):

how ?

OpenStudy (anonymous):

i dont understand this

Nnesha (nnesha):

do you want ot understand or not ? or just need an answer ?

OpenStudy (anonymous):

i need to get finished

Nnesha (nnesha):

http://www.wolframalpha.com/ then use calcualtor.

Nnesha (nnesha):

calculator*

OpenStudy (anonymous):

i did it wont work

Nnesha (nnesha):

take a screenshot how did you type that into the calc? did you put the parentheses < very important!

OpenStudy (anonymous):

just f*ck!ng tell me

Nnesha (nnesha):

1) tell me if you want to understand combination 2nd) how did you put that into the calc ?? patience watch your language plz at least you can type into the calculator. thanks.